### v3

# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

Abstract
[…]
1 Introduction
Large Language Models (LLMs) based on the Transformer architecture [attention2017] have revolutionized natural language processing. However, their immense scale, reaching hundreds of billions or trillions of parameters  [efficient2021, scaling2023], leads to prohibitive computational and memory costs, particularly GPU VRAM, hindering their widespread deployment on consumer or edge devices [edge2024, instinfer2024, dovetail2024].
Model quantization, which reduces numerical precision, is a key technique to address these bottlenecks [billm2024]. While post-training quantization (PTQ) methods like GPTQ [gptq2023] and AWQ [awq2023] compress to $\sim$4 bits, further compression (e.g., to 1-bit) typically requires quantization-aware training (QAT) [qat2018, qakd2019] to maintain performance. Indeed, QAT has demonstrated the ability to enable substantial 1-bit compression while preserving model fidelity [bitnet2023, onebit2024, fbillm2024].
However, even 1-bit models (e.g., $\sim$15.4GB for 70B parameters [billm2024]) can exceed the capacity of highly resource-constrained devices [reviewEdgeLLM2024], motivating the exploration of sub-1-bit quantization. Although prior work [stbllm2024] has ventured into this area, maintaining performance at extremely low effective bits (e.g., 0.55 bits per weight (BPW)) while ensuring computational efficiency without restrictive hardware dependencies remains a critical challenge [stbllm2024]. Consequently, it is crucial to pioneer new techniques that can achieve such profound model compression, aiming for bit rates around 0.1 BPW, while maintaining the integrity of model performance.
Figure 1: Low-bit quantization perplexity on Llama2-13B (WikiText-2). LittleBit surpasses the state-of-the-art sub-1-bit quantization technique. Below 0.5 BPW, where the leading prior method degrades sharply, ours remains robust down to 0.1 BPW.
Our approach stems from two observations. First, LLM weight matrices often exhibit significant low-rank structure [galore2024, dobi2025]. This suggests that factorization methods like Singular Value Decomposition (SVD) [svd1970] could offer a more stable compression pathway than pruning [deepcompression2016], especially under extreme compression regimes. For instance, at compression ratios exceeding 50%, SVD-based approaches [svdllmv2_2025] consistently outperform pruning methods like Wanda [wanda2023] (see Section˜A.5 for details). Second, binarization inherently causes information loss [survey2024]. Recent high-performing 1-bit methods [onebit2024, binarymos2024, arbllm2025] demonstrate that sophisticated scaling over multiple dimensions (e.g., row and column), is crucial for mitigating this loss and stabilizing performance [arbllm2025].
Based on these insights, we introduce LittleBit, a novel method for extreme sub-1-bit quantization (e.g., 0.1 BPW). LittleBit first represents weights via low-rank factorization ($\mathbf{W}\approx\mathbf{U}\mathbf{V}^{\top}$) and then binarizes these factors. To counteract errors from binarization, LittleBit leverages a multi-scale compensation mechanism that applies learnable scales across rows, columns, and an additional latent dimension. This entire architecture is then complemented by Residual Compensation.
Extensive experiments show LittleBit’s superiority over STBLLM [stbllm2024], the leading sub-1-bit technique, on LLMs from 1.3B to 32B parameters. Notably, LittleBit achieves competitive performance on Llama2-13B at an unprecedented 0.1 BPW (Fig.˜1), surpassing STBLLM at 0.55 BPW. Furthermore, the 32B LittleBit model at just 0.3 BPW maintains strong performance, significantly outperforming STBLLM and setting a new state-of-the-art for sub-1-bit quantization.
In summary, our contribution is as follows:
  • We propose LittleBit, a novel method unifying latent matrix factorization with multi-scale compensation for extreme sub-1-bit quantization. Our method achieves unprecedentedly low effective bits (e.g., down to 0.1 BPW) while preserving model performance, establishing a new, viable size-performance trade-off that provides a practical path for deploying LLMs in highly resource-constrained environments.
  • We introduce Dual-SVID for effective initialization of factorized structures and Residual Compensation to mitigate errors in extreme quantization.
  • We provide extensive validation showing that LittleBit significantly outperforms STBLLM [stbllm2024], the prior leading sub-1-bit quantization technique, on various LLM scales.
2 Related Works
2.1 Binarization and Quantization
Network binarization targets extreme compression and acceleration by constraining weights or activations to $\pm 1$, enabling efficient bitwise operations [binarized2016, towards2018]. Although early works like BinaryConnect [binaryconnect2015] and BNNs [binarized2016] showed feasibility, they faced significant accuracy degradation [binarysurvey2020]. This degradation was partially mitigated by introducing scaling factors [xnor2016] and employing the Straight-Through Estimator (STE) [ste2013] to handle the non-differentiable quantization function during training [custom2024].
However, directly applying these early techniques to sensitive large language models (LLMs) typically results in unacceptable performance loss [pbllm2023]. To address this challenge, LLM-specific quantization strategies have been developed, mainly categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [survey2021]. PTQ adapts pre-trained models, achieving competitive results around 4-bit precision with methods such as GPTQ [gptq2023] and AWQ [awq2023], but generally struggles to maintain performance below 2-bit precision [pbllm2023, billm2024]. The sub-1-bit regime is particularly challenging for PTQ methods due to severe information loss [predicting2013]. Recent work such as STBLLM [stbllm2024] shows that incorporating structured binarized compression can push PTQ into the sub-1-bit regime, yet a non-negligible accuracy gap remains.
In contrast, QAT integrates the quantization process directly into the training loop. This enables the model to learn and adapt to the low-precision format during training, generally yielding superior performance compared to PTQ at very low bit-widths [llmqat2023]. Successful QAT approaches for low-bit scenarios often employ sophisticated scaling mechanisms alongside the adaptive training itself, as evidenced by several recent methods [onebit2024, binarymos2024]. However, despite these advancements, consistently achieving high model fidelity when quantizing to less than 1.0 BPW remains a considerable challenge. This persisting difficulty underscores the need for novel architectures and training strategies specifically designed for such extreme compression regimes.
2.2 Low-Rank Approximation in Quantization
Beyond quantization, the inherent parameter redundancy within deep neural networks [predicting2013], particularly evident in LLMs [shortgpt2024, whatmatters2024], can be effectively exploited using low-rank approximation methods. Techniques such as Singular Value Decomposition (SVD) [svd1970] allow large weight matrices to be represented as products of smaller factors, offering an alternative compression strategy that can be more resilient than pruning, especially under high compression regimes [deepcompression2016]. Recognizing their potential for complementary benefits, researchers have explored combining low-rank factorization to reduce parameter count and quantization to lower parameter precision [caldera2024]. Initial studies applied SVD in various roles within the LLM context: some used activation statistics to guide decomposition [svdllmv2_2025], others integrated SVD principles into initialization or parameter-efficient updates [adalora2023, pissa2024, onebit2024], applied transformations prior to factorization [asvd2023], or even employed low-rank structures to model quantization errors [lqer2024].
Figure 2: Comparison of a standard Transformer layer (left) and the LittleBit architecture (right). LittleBit performs linear transformation using parallel Primary and Residual pathways. The Primary path employs binarized factors ($\mathbf{U}_{\mathrm{sign}},\mathbf{V}_{\mathrm{sign}}$) and FP16 scales ($\mathbf{h},\mathbf{g},\bm{\ell}$) on input $\mathbf{X}$, initialized from $\mathbf{W}$ via Dual-SVID. Simultaneously, the Residual path computes a correction with its own parameters ($\mathbf{U}_{\mathrm{res,sign}},\mathbf{V}_{\mathrm{res,sign}},\mathbf{h}_{\mathrm{res}},\mathbf{g}_{\mathrm{res}},\bm{\ell}_{\mathrm{res}}$) from the approximation residual. Their outputs sum to form $\mathbf{Y}$, eliminating storage of the effective weight matrices $\widehat{\mathbf{W}}_{\mathrm{pri}}$ and $\widehat{\mathbf{W}}_{\mathrm{res}}$.
A more integrated approach involves incorporating low-rank concepts directly within the QAT methods, as demonstrated by DL-QAT [dlqat2024], which jointly learns a representation amenable to both low-rank approximation and low-precision quantization, albeit limited to 3-bit quantization. This direction of optimizing structural properties such as rank and numerical precision during training represents a promising path towards more effective LLM compression, with potential for further exploration of even lower precision quantization.
3 Methodology
This section details LittleBit. We describe its factorized architecture with multi-scale compensation (Section˜3.1), the Dual-SVID initialization for stable QAT (Section˜3.2), and Residual Compensation for enhanced fidelity (Section˜3.3). These components are optimized via QAT (details in Section˜4).
3.1 LittleBit Architecture
The LittleBit architecture, depicted in Fig.˜2, redesigns linear layers for extreme compression (e.g., 0.1 BPW) by synergistically combining low-rank factorization and binarization. LittleBit first leverages the observed low-rank structure in LLM weights [svdllmv1_2025] to approximate the weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ as:
$\mathbf{W}\approx\mathbf{U}\mathbf{V}^{\top}\quad(\mathbf{U}\in\mathbb{R}^{d_{\mathrm{out}}\times r},\mathbf{V}\in\mathbb{R}^{d_{\mathrm{in}}\times r},r\ll\min(d_{\mathrm{in}},d_{\mathrm{out}}))$ (1)
This factorization-based approach is chosen over other structural compression methods, such as pruning, due to its superior robustness in extreme compression regimes. As demonstrated in our comparative analysis against methods like Wanda (Section˜A.5), SVD-based compression maintains model integrity far more effectively at high compression ratios, providing a more stable foundation for our ultra low-bit quantization. The resulting factors $\mathbf{U}$ and $\mathbf{V}$ are then binarized:
$\mathbf{U}_{\mathrm{sign}}=\mathrm{sign}(\mathbf{U}),\quad\mathbf{V}_{\mathrm{sign}}=\mathrm{sign}(\mathbf{V})\in\{-1,+1\}$ (2)
To compensate for the significant information loss incurred by binarization [survey2024], LittleBit incorporates learnable FP16 scales. Beyond standard row ($\mathbf{h}$) and column ($\mathbf{g}$) scaling [pbllm2023, onebit2024, arbllm2025], it introduces an additional latent scale ($\bm{\ell}\in\mathbb{R}^{r}$). This latent scale is designed to address the factorization structure by learning the relative importance of each of the $r$ latent dimensions, which correspond to the columns of the binarized factors $\mathbf{U}_{\mathrm{sign}}$ and $\mathbf{V}_{\mathrm{sign}}$:
$\mathbf{h}\in\mathbb{R}^{d_{\mathrm{out}}},\quad\mathbf{g}\in\mathbb{R}^{d_{\mathrm{in}}},\quad\bm{\ell}\in\mathbb{R}^{r}$ (3)
The primary effective weight, $\widehat{\mathbf{W}}_{\mathrm{pri}}$, is implicitly constructed from these components:
$\widehat{\mathbf{W}}_{\mathrm{pri}}=\mathrm{diag}(\mathbf{h})\mathbf{U}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{V}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{g})$ (4)
Instead of storing or optimizing $\widehat{\mathbf{W}}_{\mathrm{pri}}$ directly, LittleBit learns latent full-precision factors $\mathbf{U}$ and $\mathbf{V}$, which are binarized during the forward pass using Eq.˜2, along with the FP16 scales $\mathbf{h},\mathbf{g},\bm{\ell}$. This factorized representation enables efficient computation during the forward pass.
Proposition 1
Let $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$ be the input matrix. The forward computation $\mathbf{Y}=\mathbf{X}{\widehat{\mathbf{W}}_{\mathrm{pri}}}^{\top}$ using the primary approximation (Eq.˜4) can be efficiently computed as:
$\mathbf{Y}=((((\mathbf{X}\odot\mathbf{g})\mathbf{V}_{\mathrm{sign}})\odot\bm{\ell})\mathbf{U}_{\mathrm{sign}}^{\top})\odot\mathbf{h}$ (5)
where $\odot$ denotes element-wise multiplication with broadcasting.
This decomposition (Eq.˜5) replaces a large high precision General Matrix Multiply (GEMM) operation with two smaller binary matrix multiplications and element-wise scaling operations, offering computational advantages and reducing memory requirements. The effective bits per weight (BPW) is determined by the storage cost of these learnable parameters relative to the original matrix size (see Section˜A.4 for more details).
3.2 Dual-SVID Initialization
Naively initialing the highly constrained LittleBit structure can lead to unstable QAT. To circumvent this, we propose Dual-SVID, an SVD-based initialization method designed to provide a starting point where the initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ closely approximates $\mathbf{W}$. Dual-SVID aims to preserve essential sign and magnitude information from the optimal low-rank SVD factors (obtained from $\mathbf{W}\approx\mathbf{U}^{\prime}{\mathbf{V}^{\prime}}^{\top}$, then truncated) and map this information to the learnable LittleBit parameters ($\mathbf{U}_{\mathrm{sign}},\mathbf{V}_{\mathrm{sign}},\mathbf{h},\mathbf{g},\bm{\ell}$). The process involves: (1) Obtaining $\mathbf{U}^{\prime},\mathbf{V}^{\prime}$ via a truncated SVD. (2) Setting initial binary factors based on signs (Eq.˜6). (3) Decomposing the magnitudes $|\mathbf{U}^{\prime}|\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $|\mathbf{V}^{\prime}|\in\mathbb{R}^{d_{\mathrm{in}}\times r}$. To initialize the scales, we perform a rank-1 SVD (or equivalent rank-1 approximation) on each of these magnitude matrices separately. For $|\mathbf{U}^{\prime}|$, its best rank-1 approximation can be written as $\mathbf{s}_{U,0}(\bm{\ell}_{u,0})^{\top}$, where $\mathbf{s}_{U,0}\in\mathbb{R}^{d_{\mathrm{out}}}$ becomes the initial row scale $\mathbf{h}_{0}$, and $\bm{\ell}_{u,0}\in\mathbb{R}^{r}$ is a component contributing to the latent scale. Similarly, for $|\mathbf{V}^{\prime}|$, its rank-1 approximation $\mathbf{s}_{V,0}(\bm{\ell}_{v,0})^{\top}$ provides the initial column scale $\mathbf{g}_{0}=\mathbf{s}_{V,0}\in\mathbb{R}^{d_{\mathrm{in}}}$ and another latent scale component $\bm{\ell}_{v,0}\in\mathbb{R}^{r}$. These steps effectively disentangle multi-dimensional magnitude variations and allow us to estimate initial row ($\mathbf{h}_{0}$), column ($\mathbf{g}_{0}$), and the final latent scale ($\bm{\ell}_{0}=\bm{\ell}_{u,0}\odot\bm{\ell}_{v,0}$ by element-wise product), as shown in (Eqs.˜7 and 8):
$\mathbf{U}_{\mathrm{sign},0}=\mathrm{sign}(\mathbf{U}^{\prime}),\quad\mathbf{V}_{\mathrm{sign},0}=\mathrm{sign}(\mathbf{V}^{\prime})$ (6)
$|\mathbf{U}^{\prime}|\approx\mathbf{h}_{0}(\bm{\ell}_{u,0})^{\top},\quad|\mathbf{V}^{\prime}|\approx\mathbf{g}_{0}(\bm{\ell}_{v,0})^{\top}$ (7)
$\bm{\ell}_{0}=\bm{\ell}_{u,0}\odot\bm{\ell}_{v,0}$ (8)
Figure 3: Visualization of Dual-SVID initialized weight components for a selected layer in Llama2-7B (Query Weight, Layer 0). Columns, from left to right, represent effective bits of 0.1, 0.3, 0.55, 0.7, 0.8, and 1.0 BPW. Rows display the primary approximation ($\widehat{\mathbf{W}}_{\mathrm{pri},0}$), the residual approximation ($\widehat{\mathbf{W}}_{\mathrm{res},0}$), and their sum ($\widehat{\mathbf{W}}_{0}=\widehat{\mathbf{W}}_{\mathrm{pri},0}+\widehat{\mathbf{W}}_{\mathrm{res},0}$). The rightmost image shows the corresponding crop of the original weight matrix ($\mathbf{W}$) for reference.
The resulting initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}=\mathrm{diag}(\mathbf{h}_{0})\mathbf{U}_{\mathrm{sign},0}\mathrm{diag}(\bm{\ell}_{0})\mathbf{V}_{\mathrm{sign},0}^{\top}\mathrm{diag}(\mathbf{g}_{0})$ preserves key structural information. The learnable parameters are initialized as follows: the latent factors $\mathbf{U}$ and $\mathbf{V}$ are initialized with the factors $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ obtained from SVD respectively, and the scales $\mathbf{h},\mathbf{g},\bm{\ell}$ are initialized to $\mathbf{h}_{0},\mathbf{g}_{0},\bm{\ell}_{0}$.
3.3 Residual Compensation
To further improve fidelity without increasing the primary path’s rank, we introduce Residual Compensation, a technique motivated by the theoretical insight that a two-stage error correction can be more effective than a single, larger approximation (Section˜A.1). Crucially, this method does not increase the model’s total parameter budget; instead, it strategically reallocates the fixed bit budget of a single, higher-rank approximation into two lower-rank paths: a primary and a residual path. The term ’residual’ primarily describes its role during initialization, where the auxiliary path is configured to model the error of the primary approximation. During QAT, both paths are optimized jointly to collectively represent the original weight. The advantage of this dual-path approach over a single path is empirically validated in our ablation studies (Section˜A.2.1).
To implement this, Residual Compensation employs a parallel auxiliary path that learns an approximation $\widehat{\mathbf{W}}_{\mathrm{res}}$ of the residual error. This auxiliary path mirrors the structure of the LittleBit path:
$\widehat{\mathbf{W}}_{\mathrm{res}}=\mathrm{diag}(\mathbf{h}_{\mathrm{res}})\mathbf{U}_{\mathrm{res,sign}}\mathrm{diag}(\bm{\ell}_{\mathrm{res}})\mathbf{V}_{\mathrm{res,sign}}^{\top}\mathrm{diag}(\mathbf{g}_{\mathrm{res}})$ (9)
The parameters of this auxiliary path ($\mathbf{U}_{\mathrm{res,sign}},\mathbf{V}_{\mathrm{res,sign}},\mathbf{h}_{\mathrm{res}},\mathbf{g}_{\mathrm{res}},\bm{\ell}_{\mathrm{res}}$) are also learnable during QAT. They are initialized using the Dual-SVID strategy applied to the initial residual error $\mathbf{W}_{\mathrm{res},0}=\mathbf{W}-\widehat{\mathbf{W}}_{\mathrm{pri},0}$. Employing a separate path allows the model to specifically learn and compensate for the potentially distinct characteristics of this residual error. The final effective weight is the sum of both approximations:
$\widehat{\mathbf{W}}=\widehat{\mathbf{W}}_{\mathrm{pri}}+\widehat{\mathbf{W}}_{\mathrm{res}}$ (10)
This targeted error correction is often advantageous for maintaining performance at extreme compression levels, as demonstrated by the ablation studies presented in Section˜A.2.1. It is important to note that all learnable parameters from both the primary and residual paths are included when calculating the final BPW (see Section˜A.4 for calculation details).
Fig.˜3 visualizes the initial state of weight components ($\widehat{\mathbf{W}}_{\mathrm{pri},0}$, $\widehat{\mathbf{W}}_{\mathrm{res},0}$, and their sum $\widehat{\mathbf{W}}_{0}$) for a Llama2-7B query weight, derived via Dual-SVID across effective bits from 0.1 to 1.0, compared against the original weight $\mathbf{W}$ (far right). Even at a low 0.1 BPW, the primary approximation $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ (top row, leftmost) captures the dominant low-rank structure of $\mathbf{W}$, albeit with considerable simplification. As effective bits increase, $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ progressively refines details and more effectively suppresses disruptive approximation noise, allowing underlying weight patterns to emerge with greater clarity. A key observation is that $\widehat{\mathbf{W}}_{0}$ (bottom row) at a modest 0.3 BPW—benefiting from the initialized residual component $\widehat{\mathbf{W}}_{\mathrm{res},0}$ (middle row)—often presents a more faithful initial representation of $\mathbf{W}$ than $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ alone even at a higher 1.0 BPW. This underscores that Residual Compensation, even at the initialization stage, is vital for capturing complexities missed by the primary low-rank approximation, thereby providing a more faithful starting point for QAT.
4 Experiments
4.1 Settings
Evaluation Setup
We evaluate LittleBit on diverse LLM families including Llama [llama2023], Llama2 [llama2_2023], Llama3 [llama3_2024], OPT [opt2022], Phi-4 [phi4_2024], and QwQ [qwq32b2025], spanning sizes from 1.3B to 32B parameters. The primary performance metric is perplexity (PPL), evaluated on the WikiText-2 [wiki2_2016] validation dataset. Additional results on the C4 [C42020] and PTB [ptb1994] datasets are provided in Section˜A.3. We also assess zero-shot accuracy on common reasoning benchmarks: WinoGrande [winogrande2021], OpenBookQA (OBQA) [OpenBookQA2018], HellaSwag [hellaswag2019], BoolQ [boolq2019], ARC-Easy (ARC-e), ARC-Challenge (ARC-c) [arc2018], and PIQA [piqa2020].
Training Details
The LittleBit model parameters, initialized using the Dual-SVID method (Section˜3.2), are optimized using QAT with knowledge distillation (KD) [distilling2015, llmqat2023, ternarybert2020]. The original pre-trained full-precision model serves as the teacher ($\mathcal{T}$) for the LittleBit student model ($\mathcal{S}$). The QAT objective combines the standard output Kullback-Leibler (KL) divergence loss $\mathcal{L}_{\mathrm{out}}$ and an intermediate layer Mean Squared Error (MSE) loss $\mathcal{L}_{\mathrm{inter}}$ to match hidden representations, weighted by an empirically determined $\lambda=10$:
$\mathcal{L}_{\mathrm{QAT}}=\mathcal{L}_{\mathrm{out}}+\lambda\mathcal{L}_{\mathrm{inter}}.$ (11)
Following the setup in [binarymos2024], the training data combine WikiText-2 and selected partitions from C4. We use 2048 token sequence length, 5 epochs, Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$[adam2014], and cosine LR decay with 2% warm-up (see details in  LABEL:appendix:hyper_parameters). For models employing Grouped-Query Attention (GQA) [gqa2023] (e.g., Llama3, Phi-4, QwQ), the latent ranks $r$ for key(K) and value (V) projection layers were specifically adjusted (see Section˜A.2.3) to maintain performance under extreme quantization. For backpropagation, SmoothSign (forward: $\mathrm{sign}(x)$; backward gradient: $\tanh(kx)$, $k=100$) is used, which generally performed better than Straight-Through Estimator (STE) [ste2013] (see Section˜A.2.2). While the concept of a proxy gradient is not new, our contribution lies in the specific application and empirical validation of using the derivative of $\tanh(100x)$ as a stable gradient for QAT in the ultra-low-bit regime, which our ablation study confirms is effective (Section˜A.2.2).
Baselines
We benchmark LittleBit across a range of effective bits from approximately 1.0 down to 0.1 bits per weight (BPW). In the challenging sub-1-bit regime, we compare against STBLLM [stbllm2024], a post-training quantization (PTQ) method employing N:M sparsity. For the $\sim$1.0 BPW setting, we compare against the 1-bit PTQ method BiLLM [billm2024] and quantization-aware training (QAT) approaches (OneBit [onebit2024], BinaryMoS [binarymos2024]). Results from standard low-bit methods like GPTQ [gptq2023] and OmniQuant [omniquant2023] are included for broader context.
Table 1: Perplexity (PPL) comparison on WikiText-2 across various LLMs and quantization methods. Lower PPL indicates better performance. BPW denotes effective bits per weight. Methods marked with use quantization-aware training (QAT). STBLLM [stbllm2024] utilizes N:M sparsity (ratio in parentheses). LittleBit demonstrates strong performance, particularly in the sub-1-bit regime.
Settings OPT Llama Llama2 Llama3 Phi-4 QwQ
Method Block Size BPW 1.3B 7B 13B 7B 13B 8B 14.7B 32B
FullPrecision - 16 14.62 5.67 5.09 5.47 4.88 6.10 6.67 6.34
OmniQuant - 2 28.82 9.75 7.83 11.20 8.25 349.27 12.09 10.19
GPTQ 128 2 119.98 39.96 15.08 52.22 23.63 1.5e3 24.96 67.16
BiLLM 128 1.1 74.07 41.95 13.95 29.00 23.71 54.29 16.95 15.4
OneBit - 1 20.27 8.21 7.37 8.36 7.41 13.09 9.92 9.86
BinaryMoS - 1 18.09 7.84 7.05 7.74 6.95 10.83 9.51 8.99
LittleBit - 1 20.33 9.03 8.17 9.08 8.18 16.30 11.28 12.08
STBLLM 128 0.80 (6:8) 52.13 15.19 9.43 13.81 11.90 30.90 12.12 12.19
STBLLM 128 0.70 (5:8) 73.42 19.52 11.47 19.17 14.00 59.83 14.57 13.78
STBLLM 128 0.55 (4:8) 123.03 38.73 16.88 30.67 27.05 241.95 21.99 18.32
STBLLM 128 0.30 (2:8) 2.3e3 1.6e3 592.06 1.8e3 893.82 1.7e5 761.05 512.01
LittleBit - 0.80 21.32 9.44 8.50 9.53 8.52 17.28 11.70 12.46
LittleBit - 0.70 21.99 9.66 8.71 9.85 8.76 18.01 12.05 13.22
LittleBit - 0.55 23.35 10.09 9.16 10.47 9.24 18.47 12.80 13.57
LittleBit - 0.30 28.30 11.50 10.33 12.00 10.48 20.34 14.71 16.48
LittleBit - 0.10 53.76 15.58 13.71 15.92 15.09 26.11 19.73 35.26
Table 2: Zero-shot accuracy (%) comparison on common sense reasoning benchmarks. Compares Full Precision (FP16) models against STBLLM [stbllm2024] and LittleBit at 0.55 and 0.3 BPW.
Models Method WinoGrande OBQA HellaSwag BoolQ ARC-e ARC-c PIQA Average
Llama2-7B FullPrecision 67.16 40.80 72.95 71.37 69.44 40.95 78.12 62.97
STBLLM (0.55) 52.80 33.00 36.94 62.17 37.33 25.34 62.46 44.29
STBLLM (0.3) 50.43 31.80 26.20 37.83 25.88 25.09 53.26 35.78
LittleBit (0.55) 51.62 34.00 44.57 61.38 44.15 26.96 68.12 47.26
LittleBit (0.3) 51.30 34.80 37.61 61.80 39.98 25.09 65.83 45.20
Llama2-13B FullPrecision 69.45 41.80 76.58 69.29 73.19 44.53 78.61 64.78
STBLLM (0.55) 55.25 31.20 35.10 62.23 42.51 27.22 61.75 45.04
STBLLM (0.3) 51.54 28.80 26.01 55.63 27.53 24.74 53.32 38.22
LittleBit (0.55) 53.04 35.60 51.06 50.58 46.09 29.10 70.78 48.03
LittleBit (0.3) 51.30 34.00 43.81 55.96 42.59 25.17 68.77 45.94
Llama3-8B FullPrecision 72.92 45.00 79.18 81.25 80.21 52.98 79.54 70.15
STBLLM (0.55) 52.17 25.80 30.61 57.16 30.35 23.72 57.56 39.62
STBLLM (0.3) 49.57 26.60 26.36 51.62 26.05 24.40 51.74 36.62
LittleBit (0.55) 50.12 30.20 36.78 57.55 46.80 22.95 66.27 44.38
LittleBit (0.3) 51.93 28.20 33.91 57.98 43.48 24.32 64.64 43.49
4.2 Main Results
Superior Perplexity in Sub-1-bit Regime
Table˜1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [stbllm2024], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
Extreme Low-BPW Stability
A key strength of LittleBit is its exceptional stability and performance in the extreme sub-0.5 BPW range, specifically at 0.3 BPW and 0.1 BPW. In this territory, methods like STBLLM exhibit severe performance degradation (e.g., Llama2-7B PPL of 1.8e3 at 0.3 BPW). In contrast, LittleBit maintains strong PPL scores, such as 12.00 for Llama2-7B at 0.3 BPW and 15.92 at an unprecedented 0.1 BPW. When configured for an effective bits around 1.0 BPW by adjusting the latent rank $r$, LittleBit delivers perplexity comparable to specialized 1-bit QAT methods like OneBit [onebit2024] (e.g., Llama2-7B: LittleBit 9.08 vs. OneBit 8.36) and PTQ methods such as BiLLM [billm2024]. While a method like BinaryMoS [binarymos2024] may exhibit slightly lower PPL in some 1-bit scenarios (e.g., Llama2-7B: BinaryMoS 7.74), this can be attributed to its use of more complex mechanisms like dynamic scaling, which represent a different set of architectural and computational trade-offs. Analyzing the trade-off by adjusting latent rank to tune effective bits from 1.0 down to 0.1 BPW reveals that while performance degradation is minimal down to  0.55 BPW, a quantization cliff appears between 0.3-0.1 BPW, positioning the 0.3-0.55 BPW range as an optimal sweet spot for balancing compression and accuracy.
Model Generalization & Zero-Shot
LittleBit’s advantages are consistently observed across a diverse set of model architectures and sizes. This includes the Llama3 family, which is often considered challenging for quantization, and other contemporary models like Phi-4. For example, Llama3-8B quantized by LittleBit to 0.55 BPW yields a PPL of 18.47, significantly outperforming STBLLM’s 241.95 (at 0.55 BPW with 4:8 sparsity).The zero-shot evaluation results presented in Table˜2 further reinforce these perplexity-based findings. For instance, Llama2-7B compressed by LittleBit to 0.55 BPW achieves a mean accuracy of 47.26%, and even at an extreme 0.3 BPW, it maintains 45.20% mean accuracy. These results are notably competitive with, and in several cases surpass, those of STBLLM at similar or even higher effective bits (e.g., STBLLM Llama2-7B at 0.55 BPW scores 44.29%). This suggests a superior preservation of the models’ intrinsic reasoning capabilities even under extreme compression, highlighting LittleBit’s efficacy in creating highly compact yet powerful LLMs.
4.3 Reasoning Performance of Extremely Compressed LLMs
Figure 4: Zero-shot accuracy (%) on 7 common sense reasoning tasks for the Phi-4 14B model. Compares LittleBit-compressed Phi-4 at 0.55, 0.3, and 0.1 BPW against STBLLM.
The capability of LittleBit to achieve sub-1-bit compression while preserving significant model performance prompts an evaluation of their reasoning abilities, especially for large-scale models under extreme compression. We investigated the Phi-4 14B model, comparing its performance when compressed by LittleBit versus STBLLM across 7 zero-shot reasoning tasks. When compressed with LittleBit to 0.55 BPW, Phi-4 achieved an average accuracy of around 52.3%. While this performance is slightly lower than STBLLM at a similar compression level (0.55 BPW, 4:8 sparsity, achieving 54.2%), LittleBit demonstrates stronger graceful degradation. At a more aggressive 0.3 BPW, LittleBit-compressed Phi-4 maintained an average accuracy of 48.7%, significantly outperforming STBLLM at 0.3 BPW (2:8 sparsity), which scored 40.3%. Even at an extreme 0.1 BPW, Phi-4 compressed with LittleBit retained a notable average accuracy of 43.6%. These results highlight LittleBit’s ability to maintain more robust reasoning capabilities compared to STBLLM as compression becomes increasingly extreme, particularly in the sub-0.5 BPW regime for the Phi-4 model.
5 Analysis
Memory Footprint Reduction
[…]
Table 3: Memory footprint comparison (GB) for Llama2 [llama2_2023] models under different quantization methods. LittleBit is evaluated at various BPWs, achieved by adjusting the latent rank $r$. Compression factors relative to FP16 are shown in parentheses.The gap between the effective BPW and the compression factor arises because quantization is applied only to the Transformer blocks, while components like the embedding layer and the lm_head remain in FP16.
Model FP16 BiLLM [billm2024] OneBit [onebit2024] LittleBit (Ours)
BPW 16 1.1 1 0.8 0.55 0.3 0.1
[…]
Figure 5: Conceptual view of KV Cache storage: the standard method (left) stores the full hidden dimension ($d_{\mathrm{model}}$), whereas LittleBit (right) caches a reduced latent dimension ($r$)
Table 4: Estimated KV cache memory reduction for Llama2-7B using LittleBit, with reduction factor $\sim d_{\mathrm{model}}/r$.
BPW KV Latent Rank KV Cache
($r$) Reduction Factor
[…]
KV Cache Compression
[…]
Latency Considerations
[…]
These results highlight LittleBit’s potential for efficient deployment with significant latency reduction. The observed gap between theoretical computational gains and practical latency stems from the fact that inference at small batch sizes is often dominated by memory access rather than raw computation. Our factorized method involves multiple memory-bound steps, and our kernel, while effective, is not as exhaustively optimized as industry-standard libraries. A detailed analysis including theoretical FLOPs, practical latency, and end-to-end decoding throughput is provided in LABEL:appendix:latency.
6 Conclusion
In this work, we proposed LittleBit, a novel method advancing LLM compression to the extreme sub-0.5 BPW regime, with viable performance at 0.1 BPW. Gains were achieved via SVD-inspired latent matrix factorization, binarization of these factors, and a multi-scale (row, column, latent) compensation mechanism. We also introduced Dual-SVID initialization for stable quantization-aware training (QAT) and integrated Residual Compensation for error mitigation. Empirical results show LittleBit significantly outperforms prior sub-1-bit techniques and maintains robust fidelity across LLMs up to 32B parameters. LittleBit thus establishes a new state-of-the-art in the size–performance trade-off, significantly expanding LLM deployment in resource-constrained environments. Building on this foundation, future work can enhance practical deployment through hardware co-design for edge devices like neural processing units (NPUs). Further gains in model fidelity can be explored by integrating advanced techniques like non-uniform bit allocation, applied either across different model layers or even within a single layer by dynamically distributing the bit budget between the primary and residual paths.
Acknowledgements
We are grateful to Dr. Heonjae Ha, Dr. SangJeong Lee, Minseop Choi, Changdong Kim, and Hyochan Chong for their insightful discussions and generous support in this work.
Appendix A Appendix
A.1 Mathematical Analysis and Proofs
This section provides mathematical details for the LittleBit computation and explores theoretical aspects related to low-rank quantization errors, offering motivation for certain design choices.
Proposition 1 (LittleBit Forward Pass Computation (Proof Detail)).
As stated in Section˜3.1 (Proposition 1), for input $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$, the primary quantized weight matrix is $\widehat{\mathbf{W}}_{\mathrm{pri}}=\mathrm{diag}(\mathbf{h})\mathbf{U}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{V}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{g})$. The forward pass $\mathbf{Y}=\mathbf{X}\widehat{\mathbf{W}}_{\mathrm{pri}}^{\top}$ can be computed as shown in Eq.˜12.
$\mathbf{Y}=((((\mathbf{X}\odot\mathbf{g})\mathbf{V}_{\mathrm{sign}})\odot\bm{\ell})\mathbf{U}_{\mathrm{sign}}^{\top})\odot\mathbf{h}.$ (12)
Terms are defined in Section˜3.1.
Proof.
Starting with $\mathbf{Y}=\mathbf{X}\widehat{\mathbf{W}}_{\mathrm{pri}}^{\top}$:
$\mathbf{Y}$ $=\mathbf{X}\left(\mathrm{diag}(\mathbf{h})\mathbf{U}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{V}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{g})\right)^{\top}$
$=\mathbf{X}\,\mathrm{diag}(\mathbf{g})\mathbf{V}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{U}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{h})$
$=((((\mathbf{X}\odot\mathbf{g})\mathbf{V}_{\mathrm{sign}})\odot\bm{\ell})\mathbf{U}_{\mathrm{sign}}^{\top})\odot\mathbf{h}.$ (13)
This yields Eq.˜12, where $\odot$ implies element-wise multiplication with broadcasting. ∎
Claim 1 (Quantization Error vs. Factor Rank).
Consider a simplified model where a rank-$r$ matrix $\mathbf{W}^{(r)}=\mathbf{U}^{(r)}(\mathbf{V}^{(r)})^{\top}$ is approximated. Let $\mathbf{U}_{\mathrm{sign}}^{(r)}=\mathrm{sign}(\mathbf{U}^{(r)})$, $\mathbf{V}_{\mathrm{sign}}^{(r)}=\mathrm{sign}(\mathbf{V}^{(r)})$. If magnitudes are crudely approximated using fixed rank-1 SVDs of $|\mathbf{U}^{(r)}|$ and $|\mathbf{V}^{(r)}|$ to form scales $\mathbf{s}_{U}^{(r)},\mathbf{s}_{V}^{(r)}$:
$\widehat{\mathbf{W}}_{r}:=\left(\mathbf{U}_{\mathrm{sign}}^{(r)}\odot(\mathbf{s}_{U}^{(r)}\mathbf{1}^{\top}_{r})\right)\left(\mathbf{V}_{\mathrm{sign}}^{(r)}\odot(\mathbf{s}_{V}^{(r)}\mathbf{1}^{\top}_{r})\right)^{\top}.$ (14)
The error $\mathcal{E}(r):=\left\|\mathbf{W}^{(r)}-\widehat{\mathbf{W}}_{r}\right\|_{F}$ might be non-decreasing with $r$.
Proof Sketch / Heuristic Argument.
As rank $r$ (and thus complexity of $\mathbf{W}^{(r)}$, $\mathbf{U}^{(r)}$, $\mathbf{V}^{(r)}$) increases, the fixed-structure rank-1 magnitude approximation (via $\mathbf{s}_{U}^{(r)},\mathbf{s}_{V}^{(r)}$) may become a relatively poorer fit for the increasingly complex actual magnitudes of $\mathbf{U}^{(r)},\mathbf{V}^{(r)}$. This degradation in scaling accuracy could lead to $\mathcal{E}(r)$ not decreasing, or even increasing, motivating more sophisticated scaling as in LittleBit. ∎
Claim 2 (Quantization Bias vs. SVD Component Structure).
Let $\mathbf{W}$’s rank-$r$ SVD approximation be $\widetilde{\mathbf{W}}_{r}=\widetilde{\mathbf{W}}_{r_{1}}+\widetilde{\mathbf{R}}_{r_{2}}$ (primary $\widetilde{\mathbf{W}}_{r_{1}}$, residual $\widetilde{\mathbf{R}}_{r_{2}}$). Consider a quantization operator $\mathcal{Q}$. The relative quantization error $\|\mathcal{Q}(\mathbf{A})-\mathbf{A}\|_{F}/\|\mathbf{A}\|_{F}$ might be larger for matrices $\mathbf{A}$ with "flatter" singular value spectra or less energy concentration (e.g., $\widetilde{\mathbf{R}}_{r_{2}}$ vs. $\widetilde{\mathbf{W}}_{r_{1}}$). This suggests that $\mathcal{Q}$ applied to $\widetilde{\mathbf{R}}_{r_{2}}$ might be less effective (i.e., higher relative quantization error) than when applied to $\widetilde{\mathbf{W}}_{r_{1}}$.
Proof Sketch.
$\widetilde{\mathbf{W}}_{r_{1}}$ has dominant singular values, while $\widetilde{\mathbf{R}}_{r_{2}}$ has smaller ones, often resulting in lower energy concentration for $\widetilde{\mathbf{R}}_{r_{2}}$. If $\mathcal{Q}$’s relative error is sensitive to this (e.g., higher for lower concentration), quantizing $\widetilde{\mathbf{W}}_{r_{1}}$ and $\widetilde{\mathbf{R}}_{r_{2}}$ separately may be better than quantizing $\widetilde{\mathbf{W}}_{r}$ jointly, as it allows adapting $\mathcal{Q}$ to their differing structures. ∎
Proposition 2 (Potential Advantage of Two‑Stage Quantization of SVD Components).
Using the setup of ˜2, let $\widehat{\mathbf{W}}_{r}^{\mathrm{single}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})$ and $\widehat{\mathbf{W}}^{\mathrm{two-stage}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})+\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})$. The two-stage error $\mathcal{E}_{\mathrm{two-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}^{\mathrm{two-stage}}\right\|_{F}$ can be less than the single-stage error $\mathcal{E}_{\mathrm{single-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}_{r}^{\mathrm{single}}\right\|_{F}$. This occurs if the quantization error of the sum is greater than the sum of (vectorial) quantization errors of parts:
$\|(\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})-\widetilde{\mathbf{W}}_{r_{1}})+(\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})-\widetilde{\mathbf{R}}_{r_{2}})\|_{F}<\|\mathcal{Q}(\widetilde{\mathbf{W}}_{r})-\widetilde{\mathbf{W}}_{r}\|_{F}.$ (15)
Proof Outline and Discussion.
Let $\mathbf{E}_{\mathrm{trunc}}=\mathbf{W}-\widetilde{\mathbf{W}}_{r}$. Errors are $\Delta_{1}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})-\widetilde{\mathbf{W}}_{r_{1}}$, $\Delta_{2}=\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})-\widetilde{\mathbf{R}}_{r_{2}}$, $\Delta_{\mathrm{sum}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})-\widetilde{\mathbf{W}}_{r}$. Then $\mathcal{E}_{\mathrm{two-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-(\Delta_{1}+\Delta_{2})\|_{F}$ and $\mathcal{E}_{\mathrm{single-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-\Delta_{\mathrm{sum}}\|_{F}$. The condition Eq.˜15 makes $\mathcal{E}_{\mathrm{two-stage}}<\mathcal{E}_{\mathrm{single-stage}}$ likely. Non-linear $\mathcal{Q}$ (e.g., with adaptive scaling) can satisfy this by better handling the distinct characteristics of $\widetilde{\mathbf{W}}_{r_{1}}$ and $\widetilde{\mathbf{R}}_{r_{2}}$. This motivates LittleBit’s separate residual handling, though its residual is learned differently. ∎
A.2 Ablation Study
A.2.1 Residual Compensation
To provide further intuition behind the benefits of Residual Compensation, we visualize the output activations of selected Transformer layers under different quantization schemes. Specifically, we extract the final outputs from Transformer layers 0, 5, 10, and 15 of an OPT-1.3B model, evaluated on two randomly sampled input sequences. For each layer and input, we display a set of three activation maps corresponding to the full-precision baseline, as well as two quantized variants—with and without residual compensation. These groupings allow us to directly observe the structural differences induced by the presence or absence of the residual compensation.
As shown in  Fig.˜7, residual compensation preserves a high‑fidelity resemblance to the full‑precision baseline across all inspected layers. In the no-residual setting, the output feature distributions exhibit noticeable deformation relative to the full-precision baseline. In contrast, applying residual compensation consistently restores structural similarity to the original distribution, preserving both directional patterns and magnitude diversity.
Table˜5 presents the PPL results for the OPT-1.3B model, comparing configurations with and without residual compensation. As indicated in the table, at the extremely low bit of 0.1 BPW, the version with residual compensation (PPL 60.011) performed worse than the version without it (PPL 48.512). This suggests that for relatively small models (1.3B), when pushing effective bits to such extreme lows, the additional complexity introduced by the residual compensation mechanism might outweigh its benefits and instead hinder performance. However, for the same OPT-1.3B model at other effective bits evaluated (ranging from 0.3 BPW to 1.0 BPW), the inclusion of residual compensation consistently resulted in better (lower) PPL. Moreover, experiments conducted with other, larger models generally showed that residual compensation tended to provide superior performance. In light of these overall advantages observed across various settings and particularly for larger models, residual compensation was adopted as a standard component in all experiments presented in this study.
Input 1   Input 2
layer-0 output Refer to caption   Refer to caption
layer-5 output Refer to caption   Refer to caption
layer-10 output Refer to caption   Refer to caption
layer-15 output Refer to caption   Refer to caption
Figure 7: Visualization of Transformer layer outputs for layers 0, 5, 10, and 15 under the LittleBit quantization scheme on OPT-1.3B. For each of two randomly sampled input sequences (annotated at the top), we show horizontally grouped triplets of activation maps per layer: full-precision baseline (left), quantized with residual compensation (center), and quantized without residual compensation (right).
Table 5: Perplexity (WikiText-2) comparison between Residual and Non-Residual for OPT-1.3B at various BPWs. Learning rate was 8e-5. Lower PPL is better.
BPW Residual PPL Non-Residual PPL Difference (Residual - Non-Residual)
1.00 20.329 21.691 -1.362
0.80 21.338 22.688 -1.350
0.70 22.001 23.313 -1.312
0.55 23.457 24.788 -1.331
0.30 28.984 29.724 -0.740
0.10 60.011 48.512 11.499
A.2.2 SmoothSign versus Straight-Through Estimator
For backpropagating through the non-differentiable $\mathrm{sign}(x)$ function, we compared the Straight-Through Estimator (STE) [ste2013] with our proposed SmoothSign technique. SmoothSign employs the $\mathrm{sign}(x)$ function for the forward pass. For the backward pass, it utilizes a smooth proxy gradient, specifically the derivative of $\tanh(kx)$ where $k=100$ (illustrated in Fig.˜8).
Table˜6 (OPT-1.3B results) shows SmoothSign yielding better PPL as effective bits decrease, with a notable advantage at 0.1 BPW. Due to its superior stability and performance in the ultra-low bit regime, SmoothSign was adopted for all QAT experiments (consistent with Section˜4.1).
(a) Forward pass: $\mathrm{sign}(x)$ and $\tanh(100x)$
(b) Backward pass: gradient $\frac{d}{dx}\tanh(100x)$
Figure 8: The SmoothSign activation technique. (a) Functions relevant to the forward pass: SmoothSign uses the $\mathrm{sign}(x)$ function (actual forward pass). The plot also shows $\tanh(100x)$, which serves as the basis for deriving the smooth proxy gradient shown in (b). (b) The smooth proxy gradient, $\frac{d}{dx}\tanh(100x)$, utilized for the backward pass.
Table 6: Perplexity (WikiText-2) comparison between STE and SmoothSign for OPT-1.3B at various effective bits (BPW). Learning rate was 8e-5. Lower PPL is better.
BPW STE PPL SmoothSign PPL Difference (SmoothSign - STE)
1.00 20.322 20.329 +0.007
0.80 21.344 21.338 -0.006
0.70 22.100 22.001 -0.099
0.55 23.528 23.457 -0.071
0.30 29.058 28.984 -0.074
0.10 60.401 60.011 -0.390
A.2.3 Grouped Query Attention
Table 7: Ablation study on adjusting the latent dimension factor ($r$) for key (K) and value (V) projection layers in Llama3-8B with GQA, targeting an overall effective bit of  0.1 BPW. Performance measured by perplexity (PPL, lower is better) on WikiText-2 and C4 validation sets, and average zero-shot accuracy (%) on common sense reasoning tasks (higher is better). The ’KV Factor’ indicates the multiplier applied to the K/V layers’ latent rank $r$ compared to the standard calculation for 0.1 BPW.
KV Factor Approx. BPW WikiText-2 PPL C4 PPL Avg. Zero-shot Acc (%)
[…]
Modern LLMs often use Grouped Query Attention (GQA) [gqa2023] or Multi-Query Attention (MQA) [fast2019], where key (K) and value (V) projection layers have smaller output dimensions than query (Q) projections. Applying LittleBit at ultra-low bits (e.g., 0.1 BPW) to these GQA/MQA models can result in an extremely small latent dimension $r$ for K/V layers (e.g., $r\approx 20$ for Llama3-8B K/V layers at 0.1 BPW overall), potentially creating an information bottleneck and hindering QAT or performance.
[…]
A.3 Perplexity on C4 and PTB
To further assess LittleBit’s generalization, we evaluated its performance on additional benchmarks: the diverse C4 web text corpus and the distinct PTB news corpus (Table˜8). Focusing on Llama2-7B/13B against STBLLM, the results on these datasets confirm the trends observed on WikiText-2. LittleBit consistently achieves lower perplexity than STBLLM at comparable sub-1-bit regimes. This performance advantage becomes more significant below 0.55 BPW, where STBLLM’s performance degrades considerably, particularly on PTB.
Notably, LittleBit maintains robust performance and stability even at the extreme 0.1 BPW level across all tested datasets. This consistent ability to maintain strong language modeling capabilities under severe compression highlights the generalizability of our approach, which integrates latent factorization, multi-scale compensation, and QAT. It reinforces LittleBit’s potential for effectively deploying capable LLMs in resource-limited environments.
Table 8: Perplexity (PPL) comparison on C4 and PTB validation sets for Llama2 models. Lower PPL indicates better performance. STBLLM [stbllm2024] utilizes N:M sparsity (ratio in parentheses).
Model Method BPW C4 PTB
Llama2-7B FullPrecision 16 6.97 37.91
STBLLM (6:8) 0.80 15.42 2.4e3
STBLLM (4:8) 0.55 30.99 527.56
STBLLM (2:8) 0.30 808.98 2.3e4
LittleBit 0.80 11.77 40.75
LittleBit 0.55 12.94 50.53
LittleBit 0.30 14.78 59.17
LittleBit 0.10 19.73 83.09
Llama2-13B FullPrecision 16 6.47 50.94
STBLLM (6:8) 0.80 12.96 165.51
STBLLM (4:8) 0.55 27.38 424.82
STBLLM (2:8) 0.30 442.49 1.7e3
LittleBit 0.80 10.60 44.87
LittleBit 0.55 11.53 53.52
LittleBit 0.30 13.07 51.66
LittleBit 0.10 18.88 87.59
A.4 Average Bits Per Weight Calculation
This section details the calculation of the average effective bits per weight (which we denote as $b$ in the following equations) for a linear layer implemented using the LittleBit architecture, including the parameters from the primary ($\widehat{\mathbf{W}}_{\mathrm{pri}}$) and residual compensation ($\widehat{\mathbf{W}}_{\mathrm{res}}$) pathways, as described in Sections˜3.1 and 3.3. The calculation demonstrates how to determine the latent rank $r$ required to achieve a target value for $b$ within the range of approximately 0.1 to 1.0.
A LittleBit linear layer replaces the original weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ with two parallel structures, each comprising:
  • Two binary sign matrices: $\mathbf{U}_{\mathrm{sign}}\in\{\pm 1\}^{d_{\mathrm{out}}\times r}$ and $\mathbf{V}_{\mathrm{sign}}\in\{\pm 1\}^{d_{\mathrm{in}}\times r}$. These require $1$ bit per element.
  • Three FP16 scaling vectors: $\mathbf{h}\in\mathbb{R}^{d_{\mathrm{out}}}$, $\mathbf{g}\in\mathbb{R}^{d_{\mathrm{in}}}$, and $\bm{\ell}\in\mathbb{R}^{r}$. These require $16$ bits per element.
Since both the primary and residual paths have this structure, the total number of bits required to store the parameters for a LittleBit layer is:
[…]
$=2\times\left((d_{\mathrm{out}}\times r\times 1)+(d_{\mathrm{in}}\times r\times 1)+(d_{\mathrm{out}}\times 16)+(d_{\mathrm{in}}\times 16)+(r\times 16)\right)$
$=2r(d_{\mathrm{out}}+d_{\mathrm{in}})+32(d_{\mathrm{out}}+d_{\mathrm{in}})+32r$
Note: If Residual Compensation is not used, the initial factor of 2 should be removed from the calculation.
The average bits per weight, denoted as $b$, is calculated by dividing the total bits by the number of parameters in the original FP16 weight matrix ($d_{\mathrm{out}}\times d_{\mathrm{in}}$):
$b=\frac{2r(d_{\mathrm{out}}+d_{\mathrm{in}})+32(d_{\mathrm{out}}+d_{\mathrm{in}})+32r}{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ (16)
To achieve a target value for $b$, we rearrange Eq.˜16 to solve for the required latent rank $r$:
$r=\frac{(b\times d_{\mathrm{out}}\times d_{\mathrm{in}})-32(d_{\mathrm{out}}+d_{\mathrm{in}})}{2(d_{\mathrm{out}}+d_{\mathrm{in}})+32}$ (17)
Since $r$ must be an integer, we typically round the calculated value to the nearest suitable integer. This chosen integer $r$ then determines the actual value of $b$ achieved, which will be very close to the target value.
  • Example 1: Linear Layer ($4,096$\times$4,096$$) Let $d_{\mathrm{out}}=$4,096$\$ and $d_{\mathrm{in}}=$4,096$\$. The original number of parameters is $4,096$\times$4,096$=$16,777,216$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$8,192$\$.
    For a target $b\approx 0.55$: Using Eq.˜17 with $b=0.55$:
    $r=\frac{(0.55\times$16,777,216$)-32($8,192$)}{2($8,192$)+32}\approx 546$
    We choose integer $r=546$. The actual value of $b$ for $r=546$, using Eq.˜16, is:
    $b=\frac{2(546)($8,192$)+32($8,192$)+32(546)}{$16,777,216$}\approx 0.5498$
  • Example 2: Linear Layer ($4,096$\times$11,008$$) Let $d_{\mathrm{out}}=$4,096$\$ and $d_{\mathrm{in}}=$11,008$\$. The original number of parameters is $4,096$\times$11,008$=$45,090,816$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$15,104$\$.
    For a target $b\approx 0.1$: Using Eq.˜17 with $b=0.1$:
    $r=\frac{(0.1\times$45,090,816$)-32($15,104$)}{2($15,104$)+32}\approx 133$
    We choose integer $r=133$. The actual value of $b$ for $r=133$, using Eq.˜16, is:
    $b=\frac{2(133)($15,104$)+32($15,104$)+32(133)}{$45,090,816$}\approx 0.0999$
By following this procedure for each linear layer in the model, we can select appropriate ranks ($r$) to achieve a desired overall value for $b$ (average bits per weight), thereby controlling the trade-off between model compression and performance. The specific ranks used to achieve the average bits per weight (BPW) figures reported in the main paper (e.g., in Tables˜1 and 3) are determined using this calculation methodology for $b$ for each layer type within the respective models.
A.5 Pruning versus SVD
To select a base compression method to achieve sub-0.5 BPW, we compared pruning (Wanda [wanda2023]) with SVD-based compression (SVD-LLMv2 [svdllmv2_2025]) on Llama-7B ( Table˜9). SVD showed superior robustness at extreme compression ratios (e.g., 25% parameter retention).
Consequently, low-rank factorization was chosen to design LittleBit due to its performance resilience and the deployment advantages of dense factorized matrices over sparse pruned models (which may require specialized hardware or overhead). Although SVD-LLMv2 itself performs well, we found that initializing LittleBit’s Dual-SVID with vanilla SVD was sufficient. The subsequent QAT process effectively recovered performance, rendering the added complexity of SVD-LLMv2 for initialization unnecessary. Accordingly, we employ the simpler vanilla SVD within our Dual-SVID initialization procedure.
Table 9: Perplexity (WikiText-2) comparison between Pruning (Wanda) and SVD-based compression (SVD-LLMv2) on Llama-7B at different parameter retention ratios. Lower perplexity is better. FP16 PPL is 5.68.
Remaining Parameter Ratio Pruning [wanda2023] SVD [svdllmv2_2025]
50% 8.7 13.6
37.5% 46.0 20.1
25% 1842.7 52.6
A.6 Analysis of Generated Samples
The below generated samples illustrate qualitative differences as the BPW of the LittleBit model decreases. Two main observations are:
  1. 1.
    Impact on Specificity, Factual Detail, and Coherence: As the effective bits decreases from 0.8 BPW towards 0.1 BPW, the generated text exhibits significant changes in detail, factual accuracy, and coherence. At 0.8 BPW, the Mona Lisa sample, while attempting to provide details (e.g., "complex facial features"), becomes verbose and repetitive, and anachronistically refers to Da Vinci and the Mona Lisa in the context of the "20th century." The Turing sample at this effective bit, however, offers a standard, factually sound definition. As the effective bit drops to 0.55 BPW, the Mona Lisa sample loses coherence significantly, expressing uncertainty about the painting’s location and bizarrely shifting to a help-seeking persona ("Could you help me find the Mona Lisa painting…"). The Turing sample at 0.55 BPW begins to broaden its definition, losing some precision but remaining somewhat related.
  2. 2.
    Degradation in Factual Accuracy and Relevance at Extreme Compression: Extreme compression, especially at 0.1 BPW, leads to severe degradation in factual accuracy and relevance, as vividly demonstrated by both samples. The Mona Lisa sample at 0.1 BPW starts with a generally acceptable statement (e.g., "iconic masterpiece") but quickly devolves into hallucination, referring to the Mona Lisa as a "statue" and associating it with "Milan’s fashion world." Similarly, the Turing sample at 0.1 BPW deviates drastically from the core definition of computer science, incorrectly defining it as "the study of the theory and methods of how the human mind and the brain work" and "the science of mind." These examples underscore that while LittleBit might maintain superficial grammatical structure at very low bits, the ability to recall and generate accurate, relevant factual information is compromised, often leading to the generation of plausible-sounding but nonsensical statements.
Sample 1: Mona Lisa
Prompt: Leonardo da Vinci painted the Mona Lisa. The Mona Lisa is
BPW Generated Output (Phi-4 Model)
[…]
BPW Generated Output (Phi-4 Model)
[…]